A particle of mass m has a speed v=α/x, where x is it’s

displacement. Find the force F(x) responsible.

F(x)=-mα2

/x3


1
Expert's answer
2021-01-19T07:09:11-0500

"v=\frac{a}{x} \\

\frac{dv}{dx}=-\frac{a}{x^2} \\

\frac{dv}{dt} = \frac{dv}{dx} \times \frac{dx}{dt} \\

= v\frac{dv}{dx} \\

a = v\frac{dv}{dx} \\

= -\frac{va}{x^2} \\

= -\frac{a}{x} \times \frac{a}{x^2} \\

= -\frac{a^2}{x^3} \\

F = ma \\

= m(-\frac{a^2}{x^3}) \\

F(x) = -\frac{ma^2}{x^3}"